OpenStudy (anonymous):

Help please ! verify the identity > cos(4x)=1-8sin^2(x) +8sin^2(x)

OpenStudy (shamil98):

With what, exactly?..

OpenStudy (anonymous):

cos(4x)=1-8sin^2(x) +8sin^2(x) < verify this identity

OpenStudy (shamil98):

cos(4x) = 1 - 8sin^2(x) + 8sin^2(x) is cos(4x) = 1

OpenStudy (anonymous):

ok .. how to verify that cos(4x) =1

OpenStudy (shamil98):

it's false.

OpenStudy (anonymous):

are you sure ,it is mi assignment !